OpenStudy (anonymous):

I will give you a medal.. The length of a rectangular room is is 5 feet more than 3 times the width. The perimeter is 90 feet. Write and equation to find the dimensions of the room

OpenStudy (texaschic101):

formula for perimater of a rectangle : P = 2L + 2W P = 90 L = 3W + 5 --- length is 5 more then 3 times width now we would just sub P = 2L + 2W 90 = 2(3W + 5) + 2W that would be your equation

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

thanks